OpenStudy (anonymous):

Help please! How to solve this In a certain county, the number of charter schools is 8 less than twice the number of alternative schools. We know that there are 36 charter schools in the county. How many alternative schools are in the county? a. 44 b. 22 c. 30

OpenStudy (mrnood):

This is simply a matter of turning worde into an equation If c = no of charter schools and a = number of alternative schools then look at the first sentence the number of charter schools is 8 less than twice the number of alternative schools write an equation c=?? using that sentence from the second sentence you know that c=36, so put that number into your equation - you can solve for a

OpenStudy (anonymous):

c = 36 - 8 like this?

OpenStudy (anonymous):

30 ? isnt it :) thanks a lot for help!!!

OpenStudy (anonymous):

this is the ans let no. of alternative schools=x then according to the Q 2x - 8 = 36 2x = 36 + 8 \[x = \frac{ 36 + 8 }{ 2 }\] \[x = \frac{ 44 }{ 2 }\] x = 22 (ANS.)

OpenStudy (anonymous):

oh ok thank you so much for help! I was wrong, lol

OpenStudy (anonymous):

u r welcome

OpenStudy (anonymous):

:)